OpenStudy (anonymous):

HELP WITH THIS QUESTION PLZ ) What are the total force and the absolute pressure on the bottom of a swimming pool 2.4 m by 1.4 m whose uniform depth is 2.0 m? total force N (downward) absolute pressure N/m2 (b) What will be the pressure against the side of the pool near the bottom? N/m2 JUST NEED ANSWER

OpenStudy (blacksteel):

I suspect you'd be better served posting this problem in the physics forum.

OpenStudy (anonymous):

omg i am struggling with this question since yesterday

OpenStudy (anonymous):

HELP lol omg im like half a sleep here waiting for an answer anyone

myininaya (myininaya):

do you have an equation?

myininaya (myininaya):

or i mean formula?

OpenStudy (anonymous):

no i dont im completely lost

OpenStudy (blacksteel):

This is the wrong forum for this problem, but I can help you solve it anyway. You have a pool measuring 2.4m by 1.4m that is uniformly 2m deep. Thus, the total amount of water in the pool is 1.4*2.4*2 = 6.72 cubic meters. Water has a mass of 1000 kg/m^3, so the total mass of water in the pool is 1000*6.72 = 6720 kg. The total force acting on the pool will be the gravitational force of the water, which is found by m*g, mass*gravitational constant, where g = 9.8 m/s^2. Then the force acting on the bottom of the pool will be 6720*9.8 = 65856 kgm/s^2 = 65856 N.

OpenStudy (anonymous):

omg thank you your right wow im so happy

OpenStudy (blacksteel):

The pressure will be the total force divided by area. The area of the bottom of the pool is, of course, 2.4m * 1.4m = 3.36 m^2. Thus the absolute pressure on the bottom of the pool is 65856 / 3.36 = 19600 N/m^2

OpenStudy (anonymous):

yeah your right i just plugged it in online

OpenStudy (anonymous):

THANK YOU SO MUCHHHHHHHHHH YAY

OpenStudy (blacksteel):

No problem!

OpenStudy (blacksteel):

(That's right kids, I do physics too =P)

OpenStudy (anonymous):

thanks bye